# Geometric representations of braid and Yang-Baxter gates

### v1

@@ Abstract
Brick-wall circuits composed of the Yang-Baxter gates are integrable. It becomes an important tool to study the quantum many-body system out of equilibrium. To put the Yang-Baxter gate on the quantum computer, it has to be decomposed into the native gates of quantum computers. It is favorable to apply the least number of native two-qubit gates to construct the Yang-Baxter gate. We study the geometric representations of all X-type braid gates and their corresponding Yang-Baxter gates via the Yang-Baxterization. We find that the braid and Yang-Baxter gates can only exist on certain edges and faces of the two-qubit tetrahedron. We identify the parameters by which the braid and Yang-Baxter gates are the Clifford gate, the matchgate, and the dual-unitary gate. The geometric representations provide the optimal decompositions of the braid and Yang-Baxter gates in terms of other two-qubit gates. We also find that the entangling powers of the Yang-Baxter gates are determined by the spectral parameters. Our results provide the necessary conditions to construct the braid and Yang-Baxter gates on quantum computers.
I Introduction
Quantum computers can efficiently solve problems that are unrealistic to be solved by classical computers [1]. Simulating many-body quantum systems is the major application of quantum computers in the current NISQ era [2]. The most popular model for quantum computation is the quantum circuit model. It is the quantum version of the classical reversible computation model. The quantum circuit model is composed of the initial state, unitary evolution realized by the single- and two-qubit gates, and the final state measurements. Arbitrary unitary evolution can be approximately constructed from the single- and two-qubit gates, known as the universal gate set [3, 4].
@@ I Introduction
The essential ingredient to solve the Yang-Baxter integrable models is the $R$ matrix, the solution of the Yang-Baxter equation. In the context of quantum computation, the unitary $R$ matrix can work as the quantum gate, called the Yang-Baxter gate [13]. In the last twenty years, the interdisciplinary studies between the quantum integrable system and the quantum computation, especially based on the quantum circuit model, were developed in different directions, in either of which the Yang-Baxter gate plays the fundamental roles.
@@
The unitary solution of parameter independent Yang-Baxter equation, also called the constant Yang-Baxter equation, gives the representation of braid group [14]. Here the Yang-Baxter gate is also called the braid gate. It can characterize the low-dimensional topology but also can be viewed as the quantum gate generating the quantum entanglement [15, 16]. Therefore, the manipulation of quantum entanglement via the braid gate can be studied from the viewpoint of topology [17, 18, 19, 20]. Then various quantum information protocols, such as quantum teleportation and entanglement swapping, can be understood from the viewpoint of topological entanglement [21, 22, 23, 24].
@@
In another research direction, the Yang-Baxter gate with the spectral parameter arises from the quantum simulation (based on the circuit model) of the integrable model. It is surprising that the simulation circuit obtained from decomposing the evolution operator into two-qubit gates, called the Trotterization, is still integrable [25, 26, 27]. The spectral parameter of the Yang-Baxter gate is interpreted as the Trotter step. Taking arbitrary values for the spectral parameter, the brick-wall circuit composed with the Yang-Baxter gates is always integrable [28]. The integrable circuit preserves many nice properties of the integrable model [29, 30, 31, 32, 33, 34]. Furthermore, the integrable circuit may suggest a deep understanding of the quantum circuit model, where the evolution realized by the two-qubit gates is always factorized [35, 36, 37]. In experiments, the integrable circuit can not only benchmark the quantum computers but also provides the testbed to study the integrable-breaking dynamics [38, 39, 40, 41].
@@
Despite the vital importance of the Yang-Baxter gate, its completed characterizations are missing. For example, which two-qubit gates can be converted into the Yang-Baxter gates through the single-qubit gates? Which Yang-Baxter gates can be more easily realized on quantum computers (from the native two-qubit gates)? How to decompose the Yang-Baxter gate with the minimal applications of the native gates of quantum computers? We aim to answer the above questions by studying the geometric representation of braid and Yang-Baxter gates. Here the geometric representation refers to mapping the two-qubit gates on a three-dimensional tetrahedron, where the two-qubit gates on the same point are locally equivalent (can be converted by the single-qubit gates) [42]. Then we give the optimal gate decompositions with respect to the number of CNOT gates of the braid and Yang-Baxter gates. In our study, we consider all the X-type braid gates and the Yang-Baxter gates obtained from braid gates through the Yang-Baxterization [43, 44, 45]. We clarify the conditions under which the braid and Yang-Baxter gates are the Clifford gates [46, 47], the matchgates [48, 49], and the dual unitary gates [50]. We calculate their entangling powers, which have been applied to the study of brick-wall circuits [51].
@@
Our paper is organized as follows. In Sec. II, we review basic theories about the two-qubit gate, as well as its geometric representation. We present our results, namely the geometric representations of braid gate and Yang-Baxter gate in Secs. III and IV respectively. Sec. V is the conclusion. Appendix A gives the proof of Theorem 2. Appendix B shows the matrix expressions of the Yang-Baxter gates obtained from the Yang-Baxterization of the braid gates.
II Tetrahedron of Two-qubit gates
II.1 Nonlocal parameters and entangling power
@@
$U_{\text{core}}(\vec{a})\\ =\exp\left(\frac{i}{2}\left(a_{1}(\sigma_{x}\otimes\sigma_{x})+a_{2}(\sigma_{y}\otimes\sigma_{y})+a_{3}(\sigma_{z}\otimes\sigma_{z})\right)\right).$ (2)
@@
to denote the nonlocal parameters of $U$. We say that two-qubit gates are locally equivalent if these two-qubit gates have the same nonlocal parameters. In other words, two-qubit gates with the same nonlocal parameters only differ by some single-qubit gates.
Quantities associated with the two-qubit gate, which are invariant under the action of single-qubit gates, are called local invariants. It is expected that the local invariants are given by $[a_{1},a_{2},a_{3}]$. It has been shown that the complete set of local invariants of the two-qubit gate is given by the spectrum of $L=Q^{T}\Lambda Q$ [53], which has the eigenvalues $\Lambda=\left\{e^{i(-a_{1}+a_{2}+a_{3})},e^{i(a_{1}-a_{2}+a_{3})},e^{i(a_{1}+a_{2}-a_{3})},e^{-i(a_{1}+a_{2}+a_{3})}\right\}$. Specifically, the characteristic polynomial of $L$ is completely determined by $\mathrm{Tr}(\Lambda)$ and $\mathrm{Tr}(\Lambda^{2})$. Therefore, the local invariants are functions of $\mathrm{Tr}(\Lambda)$ and $\mathrm{Tr}(\Lambda^{2})$.
@@
$E(|\Psi\rangle)=1-\mathrm{Tr}_{1}\rho^{2},$ (8)
and $\rho=\mathrm{Tr}_{2}|\Psi\rangle\langle\Psi|$. The overline means taking the average over all product states $|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ with the uniform distribution. We call $e_{p}(U)$ as the entangling power of two-qubit gates. Because of the simple decomposition given in Eq. (1), one can find that the entangling power of the two-qubit gate is [56]
$e_{p}(U)=\frac{2}{9}\left(1-\frac{1}{16}\left|\mathrm{Tr}(\Lambda)\right|^{2}\right),$ (9)
@@
$\left|\mathrm{Tr}(\Lambda)\right|^{2}\\ =16\left(\cos^{2}a_{1}\cos^{2}a_{2}\cos^{2}a_{3}+\sin^{2}a_{1}\sin^{2}a_{2}\sin^{2}a_{3}\right).$ (10)
@@
II.2 Decomposition of two-qubit gate
@@
Any two-qubit gate $[U]=[a_{1},a_{2},a_{3}]$ can be constructed from minimal $n$ ($n\geq 3$) applications of $U^{\prime}(\theta)$ with $[U(\theta)^{\prime}]=[\theta,0,0]$, if the nonlocal parameters satisfy $0\leq a_{1}+a_{2}+a_{3}\leq n\theta$ or $a_{1}-a_{2}-a_{3}\geq\pi-n\theta$.
@@
with the single-qubit rotation gate $R_{z}(\theta)=e^{-\frac{i}{2}\theta\sigma_{z}}$. Therefore any two-qubit gate can be decomposed with six CNOTs. After some simplifications, we have the optimal CNOT decomposition for any two-qubit gate (given by the universal gate set $\{R_{z}(\theta),H,\text{CNOT}\}$) [64]
@@
with the Hadamard gate $H=(\sigma_{x}+\sigma_{z})/\sqrt{2}$ and the phase gate $S=\text{diag}\{1,i\}$. Note that we have $S\cong R_{z}(\pi/2)$. The symbol $\cong$ represents the equivalent relation up to the global phase difference. The nonlocal parameters are interchangeable because of the commutative of $\sigma_{x}\otimes\sigma_{x}$, $\sigma_{y}\otimes\sigma_{y}$, and $\sigma_{z}\otimes\sigma_{z}$. Note that the above decomposition is not optimal if $a_{3}=0$, since only two CNOTs are required.
III Geometric representations and decompositions of Braid gates
III.1 Two-qubit braid gates
@@
All solutions of $B$ in the two-qubit cases are known [65, 66]. Note that one has to add the unitary condition (to be a valid quantum gate). The core of the two-qubit gate $U_{\text{core}}$ defined in Eq. (2) is the X-type two-qubit gate. There are in total four X-type braid gates [68]. At the two-qubit basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, the four braid gates have the forms
@@
III.2 Geometric representations and decompositions
@@
We can see that the braid gates $B_{I,II,III}$ are all on the edge $A_{2}A_{3}$ of the two-qubit tetrahedron. See Fig. 3. When $B_{I,II,III}$ are not at the point $A_{2}$, namely without any vanishing nonlocal parameters, their constructions require minimal three CNOTs, therefore the decompositions in Fig. 2 are optimal. The braid gate $B_{IV}$ is locally equivalent to the CNOT.
@@
When Braid gates $B_{I,II,III}$ (with $\phi_{I,3}=0$, or $\phi_{II,2}=0$, or $\varphi_{1}=0$) are locally equivalent to the SWAP gate, they have zero entangling power.
Clifford gates are elements of the Clifford group, which preserves the Pauli group, therefore circuits only composed with Clifford gates can be classically simulated [46, 47]. Another equivalent definition of the Clifford gate is the quantum gate can be decomposed with $\{H,S,\text{CNOT}\}$. Based on the gate decomposition of the four braid gates, we know that the braid gate $B_{I}$ is a Clifford gate if $\phi_{I,1},\phi_{I,2},\phi_{I,3}\in\{k\pi/2,k\in\mathbb{Z}\}$; the braid gate $B_{II}$ is a Clifford gate if $\phi_{II,1},\phi_{II,2}\in\{k\pi/2,k\in\mathbb{Z}\}$; the braid gate $B_{III}$ is a Clifford gate if $\varphi_{1}\in\{k\pi/4,k\in\mathbb{Z}\}$ and $\varphi_{2}\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$; the braid gate $B_{IV}$ is a Clifford gate if $\varphi_{1}\in\{k\pi,k\in\mathbb{Z}\}$. We summary the results in Table 1.
Matchgate is a special family of X-type two-qubit gates. Quantum circuits only composed with the matchgates acting on the nearest neighbor qubits is another family of circuits that is classically tractable [48, 49]. Matchgate circuit can be mapped to the dynamics of the free fermion model [69]. The necessary condition of being a matchgate is to have a vanishing nonlocal parameter [70]. We can check that the braid gate $B_{I}$ is a matchgate if $\phi_{I,3}=\pi/2$; the braid gate $B_{II}$ is a matchgate if $\phi_{II,2}=\pi/2$; the braid gate $B_{III}$ can not be a matchgate; the braid gate $B_{IV}$ is always a matchgate.
It is interesting to see that two-qubit gates on the edge $A_{2}A_{3}$ belong to the dual-unitary gates, which are still unitary if the space and time indexes are exchanged [50]. The dynamical correlation functions of the brick-wall circuit composed with the dual-unitary gate can be exactly calculated. From Fig. 3, we see that braid gates $B_{I}$, $B_{II}$, and $B_{IV}$ are all dual-unitary, while $B_{IV}$ is not.
IV Geometric representations and decompositions of Yang-Baxter gates
IV.1 Yang-Baxterization

### v2

@@ Abstract
Brick-wall circuits composed of the Yang-Baxter gates are integrable. It becomes an important tool to study the quantum many-body system out of equilibrium. To put the Yang-Baxter gate on quantum computers, it has to be decomposed into the native gates of quantum computers. It is favorable to apply the least number of native two-qubit gates to construct the Yang-Baxter gate. We study the geometric representations of all X-type braid gates and their corresponding Yang-Baxter gates via the Yang-Baxterization. We find that the braid and Yang-Baxter gates can only exist on certain edges and faces of the two-qubit tetrahedron. We identify the parameters by which the braid and Yang-Baxter gates are the Clifford gate, the matchgate, and the dual-unitary gate. The geometric representations provide the optimal decompositions of the braid and Yang-Baxter gates in terms of other two-qubit gates. We also find that the entangling powers of the Yang-Baxter gates are determined by the spectral parameters. Our results provide the necessary conditions to construct the braid and Yang-Baxter gates on quantum computers.
  • June 2024
Keywords: Quantum gate, braid gate, Integrable system, Yang-Baxter equation
1 Introduction
Quantum computers can efficiently solve problems that are unrealistic to be solved by classical computers [1]. Simulating many-body quantum systems is a major application of quantum computers in the current NISQ era [2]. The most popular model for quantum computation is the quantum circuit model. It is the quantum version of the classical reversible computation model. The quantum circuit model is composed of the initial state, unitary evolution realized by certain single- and two-qubit gates, and the final state measurements. Arbitrary unitary evolution can be approximately constructed from certain single- and two-qubit gates, known as the universal gate set [3, 4].
@@ 1 Introduction
The essential ingredient to solve the Yang-Baxter integrable models is the $R$ matrix, the solution of the Yang-Baxter equation. In the context of quantum computation, the unitary $R$ matrix can work as a quantum gate, called the Yang-Baxter gate [13]. In the last twenty years, the interdisciplinary studies between the quantum integrable system and quantum computation, especially based on the quantum circuit model, were developed in different directions, in either of which the Yang-Baxter gate plays fundamental roles.
@@
Unitary solutions of parameter independent Yang-Baxter equation, also called the constant Yang-Baxter equation, give representations of braid group [14]. Here the Yang-Baxter gate is also called the braid gate. It can characterize the low-dimensional topology but also can be viewed as a quantum gate generating the quantum entanglement [15, 16]. Therefore, the manipulation of quantum entanglement via the braid gate can be studied from the viewpoint of topology [17, 18, 19, 20]. Then various quantum information protocols, such as quantum teleportation and entanglement swapping, can be understood from the viewpoint of topological entanglement [21, 22, 23, 24].
@@
In another research direction, the Yang-Baxter gate with the spectral parameter arises from the quantum simulation (based on the circuit model) of integrable models. It is surprising that the simulation circuit obtained from decomposing the evolution operator into two-qubit gates, called the Trotterization, is still integrable [25]. The spectral parameter of the Yang-Baxter gate is interpreted as the Trotter step. Taking arbitrary values of the spectral parameter, the brick-wall circuit composed of the Yang-Baxter gates is always integrable [26]. Besides the Trotterization method, the integrable circuits are also developed from the cellular automata theory, where the medium-range integrable circuits are also constructed [27, 28, 29, 30]. The integrable circuit preserves many nice properties of the integrable model. For example, the spectrum has multi-particle bound state solutions [31]; the circuit has an infinite number of conserved charges [32]; the correlation functions can be exactly computed [33, 34]; the circuit has a ballistic spin transport [35, 36]; the circuit exists strong zero modes at the boundary [37]. In experiments, the integrable circuit can not only benchmark quantum computers but also provides a testbed to study the integrable-breaking dynamics [38, 39, 40, 41].
The Yang-Baxter equation is also called the factorization equation in some literature, which describes the consistent relation for decomposing the three-body scattering matrix into the two-body ones. In the context of the quantum circuit model, the quantum computation is realized by specific single- and two-qubit gates, namely the universal gate set. Therefore, we may view the quantum circuit as a “factorized” evolution. In the sense of factorization, the integrable circuit may suggest a deep understanding of the quantum circuit model [42, 43, 44].
@@
Despite the vital importance of the Yang-Baxter gate, its completed characterizations are missing. For example, which two-qubit gates can be converted into the Yang-Baxter gates through single-qubit gates? Which Yang-Baxter gates can be more easily realized on quantum computers (from the native two-qubit gates)? How to decompose the Yang-Baxter gate with the minimal applications of the native gates of quantum computers? We aim to answer the above questions by studying the geometric representation of braid and Yang-Baxter gates. Here the geometric representation refers to mapping the two-qubit gates on a three-dimensional tetrahedron, where the two-qubit gates on the same point are locally equivalent (can be converted by single-qubit gates) [45]. Then we give the optimal decompositions (with respect to the number of CNOT) of the braid and Yang-Baxter gates. In our study, we consider all the X-type braid gates and the Yang-Baxter gates obtained from braid gates through the Yang-Baxterization [46, 47, 48]. There are other special types of two-qubit gates, such as the Clifford gate [49, 50], the matchgate [51, 52], and the dual unitary gate [53]. The circuits composed of the Clifford gate or the matchgate (acting on the nearest-neighbor qubits) can be classically simulated. The correlation function of the brick-wall circuits composed of the dual unitary gate can be exactly computed. We clarify the conditions under which the braid and Yang-Baxter gates are the Clifford gates, the matchgates, and the dual unitary gates. We calculate their entangling powers, which have been applied to the study of brick-wall circuits [54].
@@
Our paper is organized as follows. In Sec. 2, we review basic theories about the two-qubit gate, as well as its geometric representation. We present our results, namely the geometric representations of the braid gate and the Yang-Baxter gate in Secs. 3 and 4 respectively. Sec. 5 is the conclusion. A gives the proof of Theorem 2. B shows the matrix expressions of the Yang-Baxter gates obtained from the Yang-Baxterization of the braid gates.
2 Tetrahedron of Two-qubit gates
2.1 Nonlocal parameters and entangling power
@@
$U_{\text{core}}(\vec{a})=\exp\left(\frac{i}{2}\left(a_{1}(\sigma_{x}\otimes\sigma_{x})+a_{2}(\sigma_{y}\otimes\sigma_{y})+a_{3}(\sigma_{z}\otimes\sigma_{z})\right)\right).$ (2)
@@
to denote the nonlocal parameters of $U$. We say that two-qubit gates are locally equivalent if these two-qubit gates have the same nonlocal parameters. In other words, two-qubit gates with the same nonlocal parameters only differ by some single-qubit gates. In the two-qubit computational basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, the core gate has the expression
$\small U_{\text{core}}(\vec{a})=\begin{pmatrix}e^{i\frac{a_{3}}{2}}\cos\left(\dfrac{a_{1}-a_{2}}{2}\right)&0&0&ie^{i\frac{a_{3}}{2}}\sin\left(\dfrac{a_{1}-a_{2}}{2}\right)\\ 0&e^{-i\frac{a_{3}}{2}}\cos\left(\dfrac{a_{1}+a_{2}}{2}\right)&ie^{-i\frac{a_{3}}{2}}\sin\left(\dfrac{a_{1}+a_{2}}{2}\right)&0\\ 0&ie^{-i\frac{a_{3}}{2}}\sin\left(\dfrac{a_{1}+a_{2}}{2}\right)&e^{-i\frac{a_{3}}{2}}\cos\left(\dfrac{a_{1}+a_{2}}{2}\right)&0\\ ie^{i\frac{a_{3}}{2}}\sin\left(\dfrac{a_{1}-a_{2}}{2}\right)&0&0&e^{i\frac{a_{3}}{2}}\cos\left(\dfrac{a_{1}-a_{2}}{2}\right)\\ \end{pmatrix}.$ (4)
The half entries are zero. It is an example of an X-type two-qubit gate since the nonzero entries form an “X” structure. Note that the $R$ matrix of the eight-vertex model also has the “X” structure.
[…]
@@
[…]
@@
[…]
@@
2.2 Decomposition of two-qubit gate
@@
Any two-qubit gate $[U]=[a_{1},a_{2},a_{3}]$ can be constructed from minimal $n$ ($n\geq 3$) applications of $U^{\prime}(\theta)$ with $[U^{\prime}(\theta)]=[\theta,0,0]$, if the nonlocal parameters satisfy $0\leq a_{1}+a_{2}+a_{3}\leq n\theta$ or $a_{1}-a_{2}-a_{3}\geq\pi-n\theta$.
@@
with the single-qubit rotation gate $R_{z}(\theta)=e^{-\frac{i}{2}\theta\sigma_{z}}$. Applying the transformation $H\sigma_{z}H=\sigma_{x}$ or $\sqrt{\sigma_{x}}^{\dagger}\sigma_{z}\sqrt{\sigma_{x}}=\sigma_{y}$ to $e^{-\frac{i}{2}\theta(\sigma_{z}\otimes\sigma_{z})}$ gives the realization of $e^{-\frac{i}{2}\theta(\sigma_{x}\otimes\sigma_{x})}$ or $e^{-\frac{i}{2}\theta(\sigma_{y}\otimes\sigma_{y})}$. Here $H$ is the Hadamard gate given by $H=(\sigma_{x}+\sigma_{z})/\sqrt{2}$. Therefore, based on Eqs. (1) and (2), we can decompose any two-qubit gate into six CNOTs. However, it is not optimal in the number of CNOTs according to Theorem 1. After some simplifications (by moving the single-qubit gates and merging the CNOTs), we have the optimal CNOT decomposition for any two-qubit gate (given by the universal gate set $\{R_{z}(\theta),H,\text{CNOT}\}$) [67]
@@
with the phase gate $S=\text{diag}\{1,i\}$. Note that we have $S\cong R_{z}(\pi/2)$. The symbol $\cong$ represents the equivalent relation up to the global phase difference. The nonlocal parameters are interchangeable because of the commutativity of $\sigma_{x}\otimes\sigma_{x}$, $\sigma_{y}\otimes\sigma_{y}$, and $\sigma_{z}\otimes\sigma_{z}$. Note that the above decomposition is not optimal if $a_{3}=0$, since only two CNOTs are required.
3 Geometric representations and decompositions of braid gates
3.1 Two-qubit braid gates
@@
All solutions of $B$ in the two-qubit cases are known [68, 13]. Note that one has to add the unitary condition (to be a valid quantum gate). The core of the two-qubit gate $U_{\text{core}}$ defined in Eq. (2) is the X-type two-qubit gate. There are in total four X-type braid gates [20]. In the two-qubit computational basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, the four braid gates have the forms
@@
3.2 Geometric representations and decompositions
@@
In the viewpoint of integrable vertex models, $B_{I}$ and $B_{II}$ can be categorized as the six-vertex model, while $B_{III}$ and $B_{IV}$ are from the eight-vertex model. However, in the context of quantum computation, the braid gates $B_{I,II,III}$ are all on the edge $A_{2}A_{3}$ of the two-qubit tetrahedron. See Fig. 3. Correspondingly, their gate decompositions are similar and require minimal three CNOTs if $B_{I,II,III}$ are not at the point $A_{2}$. See Fig. 2. The braid gate $B_{IV}$ is a special case that is equivalent to the CNOT.
@@
Braid gates $B_{I,II,III}$ have the maximal entangling powers when $\phi_{I,3}=\pi/2$, or $\phi_{II,2}=\pi/2$, or $\varphi_{1}=\pi/4$, which are locally equivalent to the iSWAP gate defined in Eq. (7). When braid gates $B_{I,II,III}$ (with $\phi_{I,3}=0$, or $\phi_{II,2}=0$, or $\varphi_{1}=0$) are locally equivalent to the SWAP gate, they have zero entangling power. Here we can see the difficulties of characterizing the quantum entanglement via the topological invariants because not all two-qubit gates are locally equivalent to a braid gate.
Clifford gates are elements of the Clifford group, which preserves the Pauli group, therefore circuits only composed of Clifford gates can be classically simulated [49, 50]. Another equivalent definition of the Clifford gate is that the quantum gate can be decomposed with $\{H,S,\text{CNOT}\}$. Based on the gate decomposition of the four braid gates, we know that the braid gate $B_{I}$ is a Clifford gate if $\phi_{I,1},\phi_{I,2},\phi_{I,3}\in\{k\pi/2,k\in\mathbb{Z}\}$; the braid gate $B_{II}$ is a Clifford gate if $\phi_{II,1},\phi_{II,2}\in\{k\pi/2,k\in\mathbb{Z}\}$; the braid gate $B_{III}$ is a Clifford gate if $\varphi_{1}\in\{k\pi/4,k\in\mathbb{Z}\}$ and $\varphi_{2}\in\{k\pi+\pi/2,k\in\mathbb{Z}\}$; the braid gate $B_{IV}$ is a Clifford gate if $\varphi_{1}\in\{k\pi,k\in\mathbb{Z}\}$. We summarize the results in Table 1.
Matchgate is a special family of X-type two-qubit gates. Quantum circuits only composed of the matchgates acting on the nearest neighbor qubits is another family of circuits that is classically tractable [51, 52]. Matchgate circuit can be mapped to the dynamics of the free fermion model [70]. The necessary condition of being a matchgate is to have a vanishing nonlocal parameter [71]. We can check that the braid gate $B_{I}$ is a matchgate if $\phi_{I,3}=\pi/2$; the braid gate $B_{II}$ is a matchgate if $\phi_{II,2}=\pi/2$; the braid gate $B_{III}$ can not be a matchgate; the braid gate $B_{IV}$ is always a matchgate.
It is interesting to see that two-qubit gates on the edge $A_{2}A_{3}$ belong to the so-called dual-unitary gate [53]. All quantum gates are unitary. The dual unitary gate satisfies an additional unitary condition $\tilde{U}\tilde{U}^{\dagger}=\tilde{U}^{\dagger}\tilde{U}=1\!\!1$, where $\tilde{U}$ comes from the index change $\langle m|\otimes\langle n|\tilde{U}|i\rangle\otimes|j\rangle=\langle j|\otimes\langle n|U|i\rangle\otimes|m\rangle$ where $|i\rangle$, $|j\rangle$, $|m\rangle$, and $|n\rangle$ are all the computational basis. The dynamical correlation functions of the brick-wall circuit composed of the dual-unitary gate can be exactly calculated. From Fig. 3, we see that braid gates $B_{I}$, $B_{II}$, and $B_{IV}$ are all dual-unitary, while $B_{IV}$ is not.
4 Geometric representations and decompositions of Yang-Baxter gates
4.1 Yang-Baxterization